A garden 7 m by 12 m will be expanded by planting a border of flowers. The border will be of the same width around the entire garden and has an area of 92 m2.

im very confused on this problem. it also needs to be identified if its quadratic equation or not. please do tell if it is.

Answer by rothauserc(4718)

A1 = area of garden including the border (w) = (7+w)*(12+w)
A2 = area of garden = 7 * 12 = 84
Area of border = (7+w)*(12+w) - 84 = 92
(7+w)*(12+w) = 176
w^2 + 19w + 84 = 176
this is a quadratic equation
w^2 + 19w -92 = 0
factor this equation
(w-4) * (w+23) = 0
w is 4 or -23
our answer has to be w > 0, so w = 4 m